OpenStudy (anonymous):

Solve 5x2 = -30x - 65. A) x= -3 (plus/minus) 2i B) x= -3 (plus/minus) 4i C) x= -6 (plus/minus) 2i D) x= -6 (plus/minus) 4i

OpenStudy (solomonzelman):

\(\large\color{blue}{5x^2 = -30x - 65 }\) \(\large\color{blue}{5x^2 +30x =- 65 }\) \(\large\color{blue}{5(x^2 +6x )=- 65 }\) \(\large\color{blue}{5(x^2 +6x +9-9)=- 65 }\) \(\large\color{blue}{5(x^2 +6x +9)-45=- 65 }\) \(\large\color{blue}{5(x^2 +6x +9)=- 20 }\) \(\large\color{blue}{(x^2 +6x +9)=- 4 }\)

OpenStudy (solomonzelman):

tell me if you need help from here? \(\large\color{blue}{x^2 +6x +9 }\) is a perfect square trinomial.